a total of $4,500 is invested in two funds that pay 4% and 5% simple intrest. the annunal intrest is $210.00. how much is invested in the fund paying 4% interest?
Answer by checkley77(12569)

.05x+.04(4,500-x)=210
.05x+180-.04x=210
.01x=210-180
.01x=30
x=30/.01
x=3,000 amount invested @ 5%.
4,500-3,000=1,500 amount invested @ 4%.
Proof:
.05*3,000+.04*1,500=210
150+60=210
210=210